A viewing window on the submarine SEAVIEW measures 250 square inches. When near the bottom of the Mariana Trench the average gauge pressure from the water is 15000 psi. What is the total outward force on the window?
Answer #1

Force = total pressure \times area

Total Pressure = (15000 + 14.7 ) psi

force = 15014.7 \times 250 = 3.754 \times 106 pound-force